### v1

@@ A.1 Generation
SAT-Menu Data. For SAT-Menu setup, we map each instance (i.e. CNF formula) to a menu selection puzzle. The goal is to select a combination of orderable and non-orderable food items in a way that satisfies everyone’s preferences. To this end, a food-item is sampled without replacement corresponding to the list of variables in the formula. Then, every clause in the formula is treated as the preferences for an individual, leading to the creation of two distinct lists for each person: “Likes,” for food items linked to positive literals, and “Dislikes,” for those associated with negated literals. This approach is exemplified in Table D.
A.2 Statistics
Our dataset is comprised of 60,000 formulas, among which 39,909 are satisfiable (SAT) and the remaining 20,091 are unsatisfiable (unSAT) instances. For the purposes of our experiments, we randomly selected approximately 6,000 samples from this pool. As illustrated in Figure 5, the distribution of formulas according to the number of variables is detailed as follows: 3,000 formulas with 3 variables, 7,500 with 4 variables, 9,000 with 5 variables, 4,500 with 6 variables, 3,000 with 7 variables, 13,500 with 8 variables, 3,000 with 9 variables, and 16,500 with 10 variables.
In the case of SAT instances, the range of clauses ($m$) varies in [5, 90], and the alpha ($\alpha=m/n$) ranges in [1.1, 11.0]. For unSAT instances, the clauses range from 9 to 110, with the alpha varying in [2.60, 11.0]. Across the entire dataset, the average number of variables ($n$) is 7.2, the average number of clauses ($m$) is 33, and the mean $\alpha$ is 4.7. This diverse dataset provides a broad spectrum for analyzing the impact of variable and clause distribution on formula satisfiability.
Appendix B SAT Solvers
SAT solvers are tools to automatically verify the satisfiability of propositional logic formulas. The Davis–Putnam–Logemann–Loveland (DPLL) algorithm [davis1962machine] is a key component of modern SAT solvers. It consists of a backtracking-based search algorithm, enriched with deductive rules and heuristics to efficiently explore the search space and determine if a formula is satisfiable. The core backtracking algorithm simply selects a literal and assigns a truth value to it. This step produces a simplified formula where all the parts made true by the assignment are removed. This step can be applied recursively until all clauses are removed, meaning that the original formula is satisfiable. Model checkers generalize SAT solvers by counting all the models that satisfy the formula.
Given the CNF in dimacs format, we employed MiniSAT v2.2 for both solution generation (for SAT-Translate) and verification (SAT-Menu, SAT-CNF).
Appendix C Additional Results
C.1 Failure Cases
(Rishi) Discuss Box D
Appendix D Full Prompts
Box 1: SAT-Menu Prompt # System Message
@@ Appendix D Full Prompts
Box 2: SAT-CNF Prompt # System Message
@@
Box 3: SAT-Translate Prompt # System Message
@@
Box 4: SAT-CNF (Lazy Solution) # System Message
@@
Appendix E Generation Hyperparameters
Figure 6: Surface plots depict phase transition characteristics, comparing the accuracy in range $[0,1]$ against # variables and $\alpha$. The setup is SAT-Menu with zero-shot prompting. Unlike solvers, the surface plot does not show steepness with increase in # variables.
Figure 7: [Left] Phase transitions characteristics for all LLMs. [Right] Comparison of unSAT accuracy on 3-SAT decision problem. GPT-4 outperforms all other models in detecting unsatisfiable problems. Both plots were generated for SAT-Menu setup with zero-shot prompting.
Figure 8: Figure shows how accuracy improves with the increase in satisfiability ratio (in log-scale) across all LLMs. The setup used is SAT-Menu with 0-shot prompting.
Figure 9: Figure compares SAT-Menu, SAT-CNF, and SAT-Translate performances across different LLMs. In general, input type (Menu vs. CNF) does not affect the performance. SAT-Translate outperforms the rest, showing the significance of augmenting LLMs with symbolic solvers. Note, that no phase transition is observed in SAT-Translate since the translation has polynomial complexity.
Figure 10: Figure compares 0-shot and 3-shot performance on SAT-Menu comparing all LLMs. It can be observed that, except GPT-4, in-context learning does not enhance performance for other LLMs.

### v2

@@ Appendix A Dataset Statistics
Appendix B LLMs as Solvers: Failure Cases
Table 2: Configuration Parameters
Parameter Value
GPT-*
temperature 1
max_tokens 4096
top_p 1
frequency_penalty 0
presence_penalty 0
Llama 2 70B
4-bit quantization True
tokenizer type Fast
max_new_tokens 2048
batch_size 20
Mixtral $8\times 7$B
4-bit quantization True
attention Flash
max_new_tokens 4096
batch_size 30
Here, we discuss some interesting failure cases we observed during our experiments. In the SAT-CNF context, it was observed that LLMs, including GPT-4, often opt to pass the task to an external SAT solver rather than solving it themselves. Additionally, when attempting to find a solution, these models tend to provide only a conceptual outline of the solution instead of a concrete answer, a tendency that becomes more pronounced with larger formulas. When prompted explicitly for a solution, GPT-4 might simplistically conclude that the problem is unsatisfiable due to its complexity, as shown in Box C. Although this reasoning is not entirely sound – as over-constrained formulas can still potentially be solvable – it appears that LLMs might be leveraging this as a statistical feature. Note, that the number of prompt tokens increases steeply from $\approx 500$ in the underconstrained region to $\approx 1200$ in the overconstrained regions.
In the SAT-Translate approach, it was observed that LLMs (with the exception of GPT-4) frequently deviated from the specific requirements of the prompt, generating solutions that did not adhere to the syntax expected by the solver. These inconsistencies in translation led to suboptimal accuracy levels within the LLM-Modulo frameworks.
It should be noted that the input to the LLM is based on fixed templates, the rules of which can be captured using regular grammar. Thus, one could write a simple parser to map the menu input to a CNF formula. However, mapping from ambiguous natural language to solver-compliant input may be non-trivial for the LLMs.
Generally, outputs from LLMs often require additional post-processing to meet specific guidelines.
Appendix C Full Prompts
Here, we provide full prompts used in our experiments for SAT-Menu (Box C), SAT-CNF (Box C), and SAT-Translate (Box C).
Box 2: SAT-Menu Prompt # System Message
@@ Appendix C Full Prompts
Box 3: SAT-CNF Prompt # System Message
@@
Box 4: SAT-Translate Prompt # System Message
@@
Box 5: SAT-CNF (Lazy Solution) # System Message
@@
Figure 7: LLM accuracy vs. satisfiability ratio. The figure shows how accuracy improves with satisfiability ratio (in log-scale) across all LLMs. The setup used is the search version of SAT-Menu with 0-shot prompting.
Figure 8: SAT Decision vs. SAT Search comparison. The figure illustrates the phase transitions in solving random 3-SAT problems, specifically comparing the decision and search variants. Although both variants fall under the same complexity class, empirical evidence shows that LLMs more readily solve the decision problem than the search problem. Among these models, only GPT-4 demonstrates the distinct Easy-Hard-Easy phase transition characteristic. The setup is SAT-CNF with 0-shot prompting. The plot was generated using a size 4 moving window on $\alpha$ values.
Figure 9: 0-shot vs. 3-shot accuracy. The figure compares 0-shot and 3-shot performance comparing all LLMs. It can be observed that, except GPT-4, in-context learning does not enhance performance for other LLMs. The setup is the search version SAT-Menu. The plot was generated using a size 4 moving window on $\alpha$ values.
Figure 10: Confusion matrices for 3-SAT Decision Problem. The figure compares the accuracy of all LLMs on the 3-SAT decision problem. It can be observed that, except GPT-4, all other LLMs struggle to correctly classify unsatisfiable instances. The cell annotations reflect classification accuracy, normalized over the true counts (column) to account for the imbalance between SAT and unSAT instances. The setup is SAT-Menu with 0-shot prompting.
Figure 11: LLM-Modulo Framework comparison. We compare the SAT Search performance of LLM-Modulo frameworks (SAT-Translate) with standalone LLM on SAT-Menu and SAT-CNF variants. Remarkably, the SAT-Translate approach (plotted in green) outperforms the rest, showing the significance of augmenting LLMs with symbolic solvers. The plot was generated using a size 4 moving window on $\alpha$ values.
Figure 12: # Generation tokens vs. $\alpha$. The figure shows how the number of generated tokens varies with the hardness of the problem. The setup used is the search version of SAT-Menu with 0-shot prompting. The plot was generated using a size 4 moving window on $\alpha$ values.